A wire is 28 feet long. It is cut into three pieces. The longest piece is twice as long as the shortest piece and the remaining piece is four feet longer than the shortest piece. Find the length of each piece
Found 2 solutions by ankor@dixie-net.com, Solver92311:
Answer by ankor@dixie-net.com(22740)

A wire is 28 feet long. It is cut into three pieces.
a + b + c = 28
The longest piece is twice as long as the shortest piece
a = 2c
and the remaining piece is four feet longer than the shortest piece.
b = c + 4
Replace a with 2c, and b with (c+4) in the first equation
2c + (c+4) + c = 28
4c = 28 - 4
4c = 24
c = 24/4
c = 6 ft is the shortest
:
Find the length of each piece
a = 2(6)
a = 12 ft is the longest
and
b = 6 + 4
b = 10 ft is the remaining piece
:
:
Check 12 + 10 + 6 = 28

Answer by Solver92311(821)



Let represent the measure of the short piece. Then the long piece is and the middle-sized piece is .



Solve for , and then calculate and .

John

My calculator said it, I believe it, that settles it

From
I > Ø